Problem 1

Suppose that a person has a utility U = 3x1 + 9x2, where x1 is the quantity of good 1 consumed, and x2 is the quantity of good 2 consumed.

Answer the following questions, and write your answers in the Answer Sheet.

  • Find the equation representing the person’s indifference curve for a level of utility represented by the constant K. (Hint: Isolate x1 on one side.)

  • Find the person’s marginal utility with respect to good 1 (MUx1 ).

  • Find the person’s marginal utility with respect to good 2 (MUx2 ).

Find the person’s marginal rate of substitution from good 1 to good 2 (MRSx1x2 ).

Indifference Curve Equation

  
Marginal utility for good 1 (Mux1)
Marginal Utility for good 2 (MUx2)
Marginal rate of substitution from good 1 to good 2 (MRSx1x2)
Answer #1

U= K = 3x1 + 9x2

K - 3x1 = 9x2

x2= k/9 + x2/3

This is the indifference curve equation.

MUx1 = 3

MUx2= 9

MRS = MUx1/MUx2= 3/9= 1/3